OpenStudy (anonymous):

Medals included! Find each percent of change. Round to the nearest tenth of a percent. State if it is increase or decrease. From 122 to 109

jimthompson5910 (jim_thompson5910):

Is this an increase or decrease?

OpenStudy (anonymous):

decrease

jimthompson5910 (jim_thompson5910):

ok so you subtract the new amount from the old amount then you divide by the old amount finally you multiply by 100

jimthompson5910 (jim_thompson5910):

Percent Decrease = [(Old - New)/Old] * 100 Percent Decrease = [(122-109)/122] * 100 Percent Decrease = ?? I'll let you finish

OpenStudy (anonymous):

wait can you sub. words with numbers? I dont understand

jimthompson5910 (jim_thompson5910):

just did

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

is it

jimthompson5910 (jim_thompson5910):

is it what?

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

93.8?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

that's way too high

OpenStudy (anonymous):

can you guide me through it?

OpenStudy (anonymous):

please?

OpenStudy (anonymous):

Superfast?

OpenStudy (anonymous):

Medal?

jimthompson5910 (jim_thompson5910):

Percent Decrease = [(Old - New)/Old] * 100 Percent Decrease = [(122-109)/122] * 100 Percent Decrease = [13/122] * 100 Percent Decrease = 0.10655737704918 * 100 Percent Decrease = 10.655737704918% So it's roughly a 10.656% decrease

OpenStudy (anonymous):

oooooooh i see

jimthompson5910 (jim_thompson5910):

notice how if you took 10.656% of 122 and you subtracted it from 122, you would get 109 roughly

jimthompson5910 (jim_thompson5910):

think of it as a price discount if that helps

OpenStudy (anonymous):

Dude thanks so much

jimthompson5910 (jim_thompson5910):

ex: original price is $100 discount price is $90 discount percentage is 10% using the same method above, so the percent decrease is 10%